OpenStudy (anonymous):

If g(x) = 3x + 5, what is g^–1(x)?

OpenStudy (lgbasallote):

y = 3x + 5 interchange x and y x = 3y + 5 solve for y

OpenStudy (anonymous):

x+5/3=y would that be it

OpenStudy (matheducatormcg):

you needed to subtract the 5.

OpenStudy (lgbasallote):

subtract 5 then divide both sides by 3 \[y = \frac{x-5}{3}\]

OpenStudy (anonymous):

oh my bad lol i made a mistake okay thank you sooo much!